# Electrocardiogram–Language Model for Few-Shot Question Answering with Meta Learning

### v1

@@ 4.1 Implementation Details
4.1.1 ECG Encoder Pretraining
We utilize self-supervised pre-training strategy of[13] (see Section 3.2 for pre-training ECG encoder using the publicly available PhysioNet 2021 Challenge datasets [25]. Each ECG recording is sampled at 500 Hz and has a duration ranging from 5 to 144 seconds. For the global contrastive learning task, we segment each recording into 5-second segments (corresponding to 2,500 samples). During pretraining, we apply random lead masking by independently masking each lead with a probability of $p=0.5$, enhancing the model’s robustness to missing or corrupted leads. The ECG encoder is trained using the Adam optimizer with a learning rate of $5\times 10^{-5}$ for 200 epochs.
4.2 Dataset Pre-Processing
We generate fixed-length segments of 5 seconds (2,500 samples) from the ECG recordings, to standardize the input length and addressing the variability in recording durations. Due to class imbalance, we exclude classes with fewer than 140 samples for Single-Verify questions, 14 samples for Single-Choose and 50 for Single-Query question type. As explained earlier, we use question types: single-verify, single-choose, single-query, and a merged dataset comprising all three types of questions (see Table 1). Meta-training dataset $D_{\text{meta-train}}$ and meta-testing dataset $D_{\text{meta-test}}$ be composed of data points $(x_{i},q_{i},a_{i})$ drawn from their respective sets of classes $C_{\text{meta-train}}$ and $C_{\text{meta-test}}$, where $C_{\text{meta-train}}\cap C_{\text{meta-test}}=\emptyset$, ensuring disjoint class sets for training and testing. For each question type, the data was split into 80% for training and 20% for testing as shown in Table 1.
@@ 4.2 Dataset Pre-Processing
N-Way K-Shot Single-Verify Single-Choose Single-Query All-Single (Combined)
@@ 4.3 Multimodal Fusion Module Architecture
The multimodal fusion module network transforms ECG features to be compatible with the decoder-only language model, we experiment with multiple mapping approaches tailored to different aspects of feature transformation and use Attention-based Mapper as a default mechanisms due to its superior performance. It utilizes the multi-head attention mechanism with 8 heads, 4 layers, and dropout rate of 0.5. Similarly, the Linear Mapper applies a linear transformation, i.e., single layer model. Furthermore, the MLP Mapper utilizes a feed-forward neural network with $3$ layers and ReLU activation with a dropout rate of 0.5 to prevent overfitting.
@@ 4.4 Training & Inference Procedures
The optimization of the meta-learning model is performed using the AdamW optimizer with 10,000 meta-training steps and 1,000 meta-testing steps. A single meta-learning step refers to an optimization after a support set (i.e., the few-shot samples) is used by the model to learn across different tasks and a query set adapt to a new task [26]. The meta-level outer learning rate is set to $5\times 10^{-4}$, while the task-level inner update learning rate is 0.05. The inner update step in meta-learning refers to the process of adapting the model’s parameters to a specific task during inner iteration based on support set [2]. The task-level inner update steps are set to the default value of 5, and the update steps for fine-tuning are also set to the default value of 15. Due to resource limitations, we train models for one epoch (which roughly takes over a duration of 1-2 days), utilizing a step size of 10,000 split across NVIDIA H100 GPUs. We keep both ECG encoder and language model frozen, unless mentioned otherwise.
@@ 4.5 Performance Evaluation
We assess the model’s performance by comparing overlap between the generated answers and the ground truth. Given that the generated sequences may vary in length from the ground truth, we compute the accuracy by aligning the generated sequence to the length of the ground truth: $\text{Accuracy}=\frac{1}{n}\sum_{i=1}^{n}I(\hat{a}_{i}=a_{i})$, where, $\hat{a}_{i}$ is the generated token at position $i$, $a_{i}$ is the ground truth token at the same position, $n$ is the length of the ground truth sequence, and $\mathbb{I}(\cdot)$ is the indicator function, which equals $1$ if the condition is true and 00 otherwise. This metric provides a direct and fair comparison despite variations in sequence lengths inherent in text generation tasks. Furthermore, we also evaluate the model performance using various natural language generation (NLG) metrics, including BLEU[27], BertScore[28], and Rouge[29] as these have been broadly utilize to evaluate the LLM generated text [30].
@@ 5 Results
Figure 4: Qualitative analysis of Gemma-2-2B and Llama-3.1-8B models on various ECG-related question answers.
@@ 5.1 Episodic Training and Comparison with Supervised Baselines
We evaluate the effectiveness of episodic training for few-shot multimodal question answering. Table 2 presents the performance of two large language models (LLMs), Gemma-2-2B [31] and Llama-3.1-8B [32], under various few-shot settings (2-way 5-shot, 2-way 10-shot, 5-way 5-shot, and 5-way 10-shot) and question types (see Table 1, Single-Verify, Single-Choose, Single-Query, and All Single question types). We compare episodic training with standard supervised learning (Baseline) for each LLM. The results demonstrate that episodic training consistently improves performance across all settings and question types, highlighting its ability to generalize to unseen queries. Furthermore, we compare our few-shot generative approach with a fully supervised classification model adapted from image captioning to ECG question answering [5] (Upper Bound), which serves as an upper-bound on the performance. This model was trained on the original ECG-QA dataset [5] and uses exact match accuracy. In contrast, our model’s accuracy is measured by the overlap between the ground truth and the generated answer (Section 4.5) and NLG metrics in Table 3 for our key models. Our results also showcase the performance improvement achieved by using a larger LLM (Llama-3.1-8B) compared to a smaller one (Gemma-2-2B).
Furthermore, Figure 4 provides a comparative analysis of Gemma-2-2B and Llama-3.1-8B on ECG-related question answering tasks. It shows example ECGs (leads II, V1, and V6) alongside representative questions from each of the three question types. For each query, we present the ground truth (GT) and the models’ responses (A), enabling a direct visual comparison of their performance. This visualization complements the quantitative results in Table 2, offering insights into the models’ reasoning processes and their ability to extract and articulate information from ECG data across varied question formats.
@@ 5.2 Impact of Model Scale
We evaluate the 5-way 5-shot setting in single-choose question few-shot performance of several large language models (LLMs) on ECG-language question answering by simply replacing corresponding LLM in our method, including Gemma-2-2B [31], Llama-3.1-8B [32], GPT-2 [33], Phi-2-2B [34], and Qwen-2-1.5B [35]. As shown in Table 4, Llama-3.1-8B consistently achieves the highest accuracy across all question types, demonstrating a substantial performance improvement. Specifically, Llama-3.1-8B exhibits a 2.2%, 14.4%, and 27.5% improvement over the best-performing 2B parameter model (Gemma-2-2B) on S-Verify, S-Choose, and S-Query, respectively, culminating in a 24.9% overall improvement (All-S). This marked improvement suggests that the increased parameter count of Llama-3.1-8B facilitates the learning of richer representations that better capture nuanced relationships between ECG data and corresponding natural language queries. We hypothesize that utilizing a even larger LLM could potentially lead to a further significant performance improvements.
While Llama-3.1-8B exhibits superior performance, its computational requirements are substantial. Within the set of 2B parameter models, Gemma-2-2B demonstrates the strongest performance, offering a compelling balance between accuracy and computational efficiency. Accordingly, we adopt Gemma-2-2B as the default model for subsequent ablation studies, ensuring a consistent and comparable evaluation of various architectural modifications and training strategies.
@@ 5.4 Generalization on Cross-Domain Dataset
We investigate the effect of cross-domain datasets on our model’s performance under the 2-way 5-shot setting. Specifically, we evaluate the model on the MIMIC-IV-ECG dataset across different question types, with PTB-XL results provided for reference, as summarized in Table 6. As MIMIC-IV-ECG dataset is rather large, we randomly select 30000 examples from its test set for evaluation to balance computational efficiency with representativeness of the dataset.
Our method demonstrates strong cross-domain capabilities, effectively working well on the MIMIC-IV-ECG dataset when meta-adaptation techniques are incorporated. With meta-adaptation, the model achieves high accuracies of 89.7% in S-Verify and 85.7% in S-Choose question type, closely aligning with the performance on the PTB-XL dataset. This highlights the effectiveness of our approach in adapting to new domains and understanding the nuances of cross-domain data.
@@ 5.6 Model’s Capability with Reduced ECG Leads
(a) Multimodal fusion mapper
Mapper Type Accuracy
Attention Based 84.5
MLP 60.9
Linear 72.5
(b) ECG encoder training
ECG Encoder Accuracy
Frozen 84.5
Unfrozen 76.7
(c) Meta-knowledge impact
Variants Accuracy
w meta knowledge 84.5
w/o meta knowledge 0.3
@@ 5.7 Architectural Components Ablation
5.7.1 Multimodal Fusion Mapper
@@
5.7.2 Freezing ECG Encoder Parameters
@@
5.7.3 Meta-Knowledge Incorporation
Incorporating meta-knowledge significantly improves performance on few-shot learning tasks. Meta-knowledge refers to information about the learning process itself, such as patterns or strategies learned from previous tasks that can be applied to new tasks with limited data. Table 9 provide these results, where, our model achieved 84.5% accuracy on single-choice questions when leveraging meta-knowledge. Without meta-knowledge, accuracy dropped drastically to 0.3%. This substantial difference highlights the critical role of meta-knowledge in enhancing the model’s ability to effectively utilize prior information for improved understanding and decision-making in few-shot scenarios.
5.7.4 Impact of Prompt Format on Model Performance
@@ 6 Conclusion
In this work, we introduce a LLM-agnostic multimodal meta-learning framework for few-shot ECG-language question answering, addressing the critical challenges of limited labeled data and evolving task distribution in ECG interpretation. Our framework seamlessly integrates ECG signals with text queries through a trainable multimodal fusion mapper. The empirical evaluation demonstrates superior generalization performance across a range of language models, diverse few-shot learning scenarios, and varying question types. These results underscore the potential of our framework to enhance clinical practice by enabling rapid adaptation to new tasks and patient populations. We contribute to the broader field of AI in healthcare by showcasing the efficacy of combining meta-learning techniques with multimodal fusion in ECG interpretation, paving the way for efficient AI-assisted diagnostic systems. While promising, our approach has limitations. The current focus on a single ECG restricts the generalizability of our findings to comparison questions involving more than one ECG examples. Future research could explore incorporating vision modality (e.g., chest X-ray images) to develop more comprehensive models.
Reference

### v2

@@ 4.1 Implementation Details
We utilize a self-supervised pre-training strategy of (Oh et al., 2022) (see Section 3.2) for pre-training ECG encoder using the publicly available PhysioNet 2021 Challenge datasets (Goldberger et al., 2000 (June 13). Each ECG recording is sampled at 500 Hz and has a duration ranging from 5 to 144 seconds. For the global contrastive learning task, we segment each recording into 5-second segments (corresponding to 2,500 samples). The rest of the implementation details are provided in Appendix A.1.
4.2 Pre-Processing
Due to class imbalance, we exclude data points of classes with fewer than 140 samples for Single-Verify questions, 14 samples for Single-Choose, and 50 for Single-Query question types as described further in Appendix A.2.
@@ 4.2 Pre-Processing
N-Way K-Shot S-Verify S-Choose S-Query All-Single (Combined)
@@ 4.3 Multimodal Fusion Module Architecture
We experiment with multiple mapping approaches tailored to different aspects of feature transformation and use Attention-based Mapper (see in Appendix A.3 ) as a default mechanism due to its superior performance.
@@ 4.4 Training & Inference Procedures
The optimization of the meta-learning model is performed using the AdamW optimizer with 10,000 meta-training steps and 1,000 meta-testing steps. Rest of the training details are provided in Appendix A.4.
@@ 4.5 Performance Evaluation
We assess the model’s performance by comparing the overlap between the generated answers and the ground truth. Given that the generated sequences may vary in length from the ground truth, we compute the accuracy by aligning the generated sequence to the length of the ground truth: $\text{Accuracy}=\frac{1}{n}\sum_{i=1}^{n}I(\hat{a}_{i}=a_{i})$, where, $\hat{a}_{i}$ is the generated token at position $i$, $a_{i}$ is the ground truth token at the same position, $n$ is the length of the ground truth sequence, and $\mathbb{I}(\cdot)$ is the indicator function, which equals $1$ if the condition is true and 00 otherwise. Furthermore, we also evaluate the model’s performance using various natural language generation (NLG) metrics, including BLEU (Papineni et al., 2002), BertScore (Zhang et al., 2019), and Rouge (Lin, 2004) as these have been broadly utilized to evaluate the LLM generated text (Abbasian et al., 2024).
@@ 5.1 Episodic Training and Comparison with Supervised Baselines
We evaluate the effectiveness of episodic training for few-shot multimodal question answering. Table 2 presents the performance of two large language models (LLMs), Gemma-2-2B (Team et al., 2024) and Llama-3.1-8B (Dubey et al., 2024), under various few-shot settings (2-way 5-shot, 2-way 10-shot, 5-way 5-shot, and 5-way 10-shot) and question types (see Table 1, Single-Verify, Single-Choose, Single-Query, and All Single question types). We compare episodic training with standard supervised learning (Baseline) for each LLM. The results demonstrate that episodic training consistently improves performance across all settings and question types, highlighting its ability to generalize to unseen queries. Furthermore, we compare our few-shot generative approach with a fully supervised classification model adapted from image captioning to ECG question answering (Oh et al., 2024) (Upper Bound), which serves as an upper-bound on the performance. This model was trained on the original ECG-QA dataset (Oh et al., 2024) and uses exact match accuracy. In contrast, our model’s accuracy is measured by the overlap between the ground truth and the generated answer (Section 4.5) and NLG metrics in Table 3 for our key models. Our results also showcase the performance improvement achieved by using a larger LLM (Llama-3.1-8B) compared to a smaller one (Gemma-2-2B).
Furthermore, Appendix B.2 Figure 3 provides a comparative analysis of Gemma-2-2B and Llama-3.1-8B on ECG-related question answering tasks. It shows example ECGs (leads II, V1, and V6) alongside representative questions from each of the three question types. For each query, we present the ground truth (GT) and the models’ responses (A), enabling a direct visual comparison of their performance. This visualization complements the quantitative results in Table 2, offering insights into the models’ reasoning processes and their ability to extract and articulate information from ECG data across varied question formats.
@@ 5.2 Impact of Model Scale
We evaluate the 5-way 5-shot setting in single-choose question few-shot performance of several large language models (LLMs) on ECG-language question answering by simply replacing the corresponding LLM in our method, including Gemma-2-2B (Team et al., 2024), Llama-3.1-8B (Dubey et al., 2024), GPT-2 (Radford et al., 2019), Phi-2-2B (Javaheripi et al., 2023), Qwen-2-1.5B (Bai et al., 2023), SmolLM-2-1.7B (Allal et al., 2025), DeepSeek-R1-1.5B (Guo et al., 2025). As shown in Table 4, Llama-3.1-8B consistently achieves the highest accuracy across all question types, demonstrating a substantial performance improvement. Specifically, Llama-3.1-8B exhibits a 2.2%, 14.4%, and 27.5% improvement over the best-performing 2B parameter model (Gemma-2-2B) on S-Verify, S-Choose, and S-Query, respectively, culminating in a 24.9% overall improvement (All-S). This marked improvement suggests that the increased parameter count of Llama-3.1-8B facilitates the learning of richer representations that better capture nuanced relationships between ECG data and corresponding natural language queries. We hypothesize that utilizing an even larger LLM could potentially lead to further significant performance improvements.
While Llama-3.1-8B exhibits superior performance, its computational requirements are substantial. Within the set of 2B parameter models, Gemma-2-2B demonstrates the strongest performance, offering a compelling balance between accuracy and computational efficiency. Accordingly, we adopt Gemma-2-2B as the default model for subsequent ablation studies.
@@ 5.4 Generalization on Cross-Domain Dataset
We investigate the effect of cross-domain datasets on our model’s performance under the 2-way 5-shot setting. Specifically, we evaluate the model on the MIMIC-IV-ECG dataset across different question types, with PTB-XL results provided for reference, as summarized in Table 6. As the MIMIC-IV-ECG dataset is rather large, we randomly select 30,000 examples from its test set for evaluation to balance computational efficiency with representativeness of the dataset.
Our method demonstrates strong cross-domain capabilities, effectively working well on the MIMIC-IV-ECG dataset when meta-adaptation techniques are incorporated. With meta-adaptation, the model achieves high accuracies of 89.7% in S-Verify and 85.7% in S-Choose question types, closely aligning with the performance on the PTB-XL dataset. This highlights the effectiveness of our approach in adapting to new domains and understanding the nuances of cross-domain data.
@@ 5.6 Model’s Capability with Reduced ECG Leads
Multimodal fusion mapper Mapper Type Accuracy Attention Based 84.5 MLP 60.9 Linear 72.5
ECG encoder training ECG Encoder Accuracy Frozen 84.5 Unfrozen 76.7
Meta-knowledge impact Variants Accuracy w meta knowledge 84.5 w/o meta knowledge 0.3
@@ 5.7 Architectural Components Ablation
Multimodal Fusion Mapper.
@@
Freezing ECG Encoder Parameters.
@@
Meta-Knowledge Incorporation.
Incorporating meta-knowledge significantly improves performance on few-shot learning tasks. Meta-knowledge refers to information about the learning process itself, such as patterns or strategies learned from previous tasks that can be applied to new tasks with limited data (Finn et al., 2017). Table 9 provides these results, where our model achieved 84.5% accuracy on single-choice questions when leveraging meta-knowledge. Accuracy dropped drastically to 0.3% without learning meta-knowledge, highlighting the critical role of prior information for improved understanding and decision-making in few-shot scenarios.(Rafiei et al., 2024)
Impact of Prompt Format on Model Performance.
@@ 6 Conclusion
In this work, we introduce a LLM-agnostic multimodal meta-learning framework for few-shot ECG-language question answering, addressing the critical challenges of limited labeled data and evolving task distribution in ECG interpretation. Our framework seamlessly integrates ECG signals with text queries through a trainable multimodal fusion mapper. The empirical evaluation demonstrates superior generalization performance across a range of language models, diverse few-shot learning scenarios, and varying question types. These results underscore the potential of our framework to enhance clinical practice by enabling rapid adaptation to new tasks and patient populations. Our method can be easily extended to multiple ECG comparisons by incorporating multiple ECG prefixes in the LLM decoder. Future research could explore incorporating vision modality (e.g., chest X-ray images) to develop more comprehensive models. Additionally, investigating different ECG encoder variants to enhance model robustness across different patient demographics, hospitals, and ECG devices. Leveraging larger language models (LLMs), and integrating more established few-shot learning methods over multiple, randomly seeds could further improve performance and generalizability.
Acknowledgments
We acknowledge the use of the Dutch National Supercomputer Snellius for essential computational tasks.
Appendix A Implementation Details
A.1 ECG Encoder Pretraining Parameters
During pretraining, we apply random lead masking by independently masking each lead with a probability of $p=0.5$, enhancing the model’s robustness to missing or corrupted leads. The ECG encoder is trained using the Adam optimizer with a learning rate of $5\times 10^{-5}$ for 200 epochs.
A.2 Dataset Pre-Processing Details
Meta-training dataset $D_{\text{meta-train}}$ and meta-testing dataset $D_{\text{meta-test}}$ are composed of data points $(x_{i},q_{i},a_{i})$ drawn from their respective sets of classes $C_{\text{meta-train}}$ and $C_{\text{meta-test}}$, where $C_{\text{meta-train}}\cap C_{\text{meta-test}}=\emptyset$, ensuring disjoint class sets for training and testing. For each question type, the data were split into 80% for training and 20% for testing.
A.3 Multimodal Fusion Module Architecture Parameters
Attention-based Mapper utilizes the multi-head attention mechanism with 8 heads, 4 layers, and a dropout rate of 0.5. Similarly, the Linear Mapper applies a linear transformation, i.e., a single-layer model. Furthermore, the MLP Mapper utilizes a feed-forward neural network with $3$ layers and ReLU activation with a dropout rate of 0.5 to prevent overfitting.
A.4 Training & Inference Procedures Parameters
The meta-level outer learning rate is set to $5\times 10^{-4}$, while the task-level inner update learning rate is 0.05. The inner update step in meta-learning refers to the process of adapting the model’s parameters to a specific task during inner iteration based on the support set (Finn et al., 2017). The task-level inner update steps are set to the default value of 5, and the update steps for fine-tuning are also set to the default value of 15. Due to resource limitations, we train models for one epoch (which roughly takes over a duration of 1-2 days), utilizing a step size of 10,000 split across NVIDIA H100 GPUs. We keep both the ECG encoder and language model frozen, unless mentioned otherwise. Implementation details like seeds will be released with our code.
Appendix B Additional Figures and Analysis
This appendix contains additional figures and analysis that provide further insights into our experiments and results. The following subsections detail class formation, attribute distribution, meta-learning processes, and qualitative analysis of ECG-related question answers.
B.1 Meta-Training and Meta-Testing Processes
Figure 2: Illustration of the meta-training and meta-testing processes of our approach.
Following the Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) structure, we train the model on a variety of ECG question-answering tasks in the meta-training phase to make it optimize the model’s ability to quickly adapt to new tasks with minimal data. We highlight how the model’s parameters are adjusted across multiple training episodes, leading to improved accuracy in the few-shot settings presented in the study. We demonstrate the key components of the process that are critical for understanding how the models adapt in figure 2.
B.2 ECG-Related Question Answering: Qualitative Analysis
Figure 3: Qualitative analysis of Gemma-2-2B and Llama-3.1-8B models across randomly selected questions.
This figure 3 presents qualitative results comparing two models, Gemma-2-2B and Llama-3.1-8B, on single-verify, single-choose, single-query, 3 ECG-related question types. The analysis helps in understanding the models’ performance and their ability to handle various question forms.